# Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback

### v1

@@ Abstract
A schematic of kinetic temperatures is shown in Fig. 2. The kinetic temperatures $T_{u}$ and $T_{v}$ correspond to scaling the Gibbs state temperature $T^{\mathrm{G}}$ according to the widths in eigen-directions of the distribution.
We next define the work rate as follows. First, the internal energy of the system is given by $E(t)=\mathrm{tr}[\hat{H}_{\mathrm{HO}}\hat{\rho}(t)]$. Then, based on the first law of thermodynamics, the work rate corresponding to the energy extracted from the system is defined as $\dot{W}_{\mathrm{ext}}:=\dot{Q}-\frac{d}{dt}E(t)$, where $\dot{Q}:=\mathrm{tr}\left(\hat{H}_{\mathrm{HO}}\mathcal{L}_{\mathrm{bath}}[\rho]\right)$ is the heat Gong et al. (2016). The time variation of the energy can be expressed as contributions along the directions of $\sigma_{uu}$ and $\sigma_{vv}$, using the quadrature operators $\hat{u}$ and $\hat{v}$ associated with these directions, as $\frac{d}{dt}E(t)=\frac{\hbar\omega}{2}\frac{d}{dt}\mathrm{tr}\left[\hat{\rho}\hat{u}^{2}\right]+\frac{\hbar\omega}{2}\frac{d}{dt}\mathrm{tr}\left[\hat{\rho}\hat{v}^{2}\right]$. Similarly, the work rate can be further decomposed as:
$\dot{W}_{\mathrm{ext}}=\dot{W}_{\mathrm{ext},u}+\dot{W}_{\mathrm{ext},v},$ (7)
where the terms $\dot{W}_{\mathrm{ext},u}$ and $\dot{W}_{\mathrm{ext},v}$ represent the contributions of the work rate along the eigen-directions of $\bm{\sigma}$.
Main results.— We now present the main result of this study, a fundamental bound of the work rate in terms of kinetic temperatures. Suppose that the system is in the steady state; the unconditional state $\hat{\rho}$ is given by the steady solution of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). Then, we have a generalized second law of thermodynamics in the form
This inequality indicates how much QC-transfer entropy can be utilized to enhance work extraction and lower the kinetic temperatures, and serves as a quantum generalization of the second-law-like bound for classical Langevin systems Horowitz and Sandberg (2014).
We compare our result (8) with the generalized second law using the temperature $T$ of the heat bath, given in the form $\dot{W}_{\rm{ext}}/T\leq\dot{I}_{\mathrm{QCT}}$ Yada et al. (2022). When feedback cooling reduces the kinetic temperatures $T_{u}$ and $T_{v}$ below the bath temperature $T$, we have $\dot{W}_{\rm{ext}}/T\leq\dot{W}_{\mathrm{ext},u}/T_{u}+\dot{W}_{\mathrm{ext},v}/T_{v}$. In such a case, inequality (8) provides a tighter bound, as it explicitly involves the kinetic temperatures of the system itself. Since cooling is achieved by extracting work from the system and lowering its kinetic temperatures, inequality (8) represents the fundamental cooling limit.
We next consider the conditions for achieving the equality in (8). A key ingredient is the quantum Kalman filter, which provides the conditional expectation values as the estimates, as presented in Eq. (3). Note that $\hat{\rho}_{t}^{e_{t}}=\hat{\rho}_{\mathrm{c}}$ holds in this case. By using this estimate based on the quantum Kalman filter, one can implement the so-called state-based feedback Doherty and Jacobs (1999); Magrini et al. (2021), where the feedback Hamiltonian (2) is given by the coefficients $c_{x}=b_{x}\langle\hat{x}\rangle_{\mathrm{c}}+b_{p}\langle\hat{p}\rangle_{\mathrm{c}}$ and $c_{p}=-a_{x}\langle\hat{x}\rangle_{\mathrm{c}}-a_{p}\langle\hat{p}\rangle_{\mathrm{c}}$. This feedback Hamiltonian aims to reduce the mean position and momentum to the origin in the phase space. Another requirement for achieving the equality in (10) is a large feedback gain. In fact, the equality in (10) is asymptotically achieved as the feedback gain goes to infinity if the estimate is implemented by the quantum Kalman filter (see the Supplemental Material sup for the proof). For simplicity, let the parameters of the feedback Hamiltonian be $a_{x}=b_{p}=g$ and $a_{p}=b_{x}=0$. In this case, the equality in (10) is asymptotically achieved in the limit $g\to\infty$. This implies that, under these conditions, the feedback system perfectly converts the acquired information into work.
Under the same conditions, the occupancy number of the harmonic Hamiltonian $\hat{H}_{\rm HO}$, $\langle a^{\dagger}a\rangle$, also reaches its minimal value determined by the bath temperature and the measurement rate (see Supplemental Material sup ), as the optimal feedback minimizes the sum of the variances of position and momentum, which is equivalent to the occupation number in our setup. This highlights the connection between the information-theoretic bound on thermodynamic energy conversion and the cooling limit of Gaussian systems.
@@
Outline of the proof, and more on quantum information flow.— To outline the derivation of our main inequality (8) and to clarify the structure behind its equality condition, here we introduce yet another information quantity, named the quantum-classical (QC) information flow, which is slightly different from the QC-transfer entropy. The QC information flow is defined as
@@
where $\chi(\rho:e)\coloneqq S(\rho)-\mathbb{E}_{e}\left[S(\rho^{e})\right]$ is the Holevo information Holevo (1973). Since the Holevo information is the measure of the correlation between a quantum state and a classical memory, the QC information flow $-\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}$ quantifies the change in the system-memory correlation induced by the time evolution of the system during the feedback. We note that the QC information flow is a quantum version of the classical information flow Horowitz and Esposito (2014), which is defined as the partial derivative of the classical mutual information.
@@
where $\hat{\rho}^{e_{t}}_{t,\mathrm{BA}}:=\hat{\rho}_{t}^{e_{t}}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{t}^{e_{t}}]dt$ is the state evolved for a time interval $dt$ under the contributions of measurement backaction. The term $-\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}-\dot{S}_{\mathrm{BA}}$ quantifies the information available for feedback control, because the unavoidable entropy change due to the measurement backaction is subtracted from the change in correlation originating from the dynamics of the system. The detailed derivation of (10) is provided in the Supplemental Material sup . A sufficient condition to achieve the equality of (10) is the quantum Kalman filter and a large feedback gain.
We next relate the two measures of quantum information flow: QC-transfer entropy and QC information flow. Again suppose that the system is in the steady state of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). Then, the QC-transfer entropy provides an upper bound of the QC information flow:
@@
See Supplemental Material sup for the derivation. A sufficient condition for the equality in (12) is the quantum Kalman filter. For the classical case of (12), the equality condition can be expressed in terms of sufficient statistics Matsumoto and Sagawa (2018).

### v2

@@ III.1 Information thermodynamic cooling limit
III.2 Numerical demonstration
We present a numerical demonstration of our main inequality (14), by numerically implementing the quantum Kalman filter. Figure 3 shows the behavior of ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ and ${\dot{W}_{\rm{ext}}}/{T}$ as functions of the feedback gain in the steady state of Eq. (II.1). The parameters are taken from Ref. Magrini et al. (2021), an experiment of feedback cooling of levitated nanoparticles.
[…]
In addition, we compared our measurement scheme with an alternative scheme that produces the same heating rate for the occupation number. We performed numerical simulations for both methods and demonstrated that our protocol has an advantage in cooling (see Sec. E of the appendix).
III.3 Outline of the proof, and more on quantum information flow
@@
To outline the derivation of our main inequality (14) and to clarify the structure behind its equality condition, here we introduce yet another information quantity, named the quantum-classical (QC) information flow, which is slightly different from the QC-transfer entropy. The QC information flow is defined as
@@
where $\chi(\rho:e)\coloneqq S(\rho)-\mathbb{E}_{e}\left[S(\rho^{e})\right]$ is the Holevo information Holevo (1973). Since the Holevo information is the measure of the correlation between a quantum state and a classical memory, the QC information flow $\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}$ quantifies the change in the system-memory correlation induced by the time evolution of the system during the feedback. We note that the QC information flow is a quantum version of the classical information flow Horowitz and Esposito (2014), which is defined as the partial derivative of the classical mutual information.
@@
where $\hat{\rho}^{e_{t}}_{t,\mathrm{BA}}:=\hat{\rho}_{t}^{e_{t}}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{t}^{e_{t}}]dt$ is the state evolved for a time interval $dt$ under the contributions of measurement backaction. The term $\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}-\dot{S}_{\mathrm{BA}}$ quantifies the information available for feedback control, because the unavoidable entropy change due to the measurement backaction is subtracted from the change in correlation originating from the dynamics of the system. The detailed derivation of (18) is provided in Sec. B.1 of the appendix. A sufficient condition to achieve the equality of (18) is the quantum Kalman filter and a large feedback gain.
We next relate the two measures of quantum information flow: QC-transfer entropy and QC information flow. Again suppose that the system is in the steady state of Eq. (II.1). Then, the QC-transfer entropy provides an upper bound of the QC information flow:
@@
See Sec. B.2 of the appendix for the derivation. A sufficient condition for the equality in (20) is the quantum Kalman filter. For the classical case of (20), the equality condition can be expressed in terms of sufficient statistics Matsumoto and Sagawa (2018).